### v2

# VibeCodeHPC: An Agent-Based Iterative Prompting Auto-Tuner for HPC Code Generation Using LLMs

Abstract
In this study, we propose VibeCodeHPC, a multi-agent system based on large language models (LLMs) for the automatic tuning of high-performance computing (HPC) programs on supercomputers. VibeCodeHPC adopts Claude Code as its backend and provides an integrated environment that facilitates program development in supercomputer settings. The system not only brings the Vibe Coding paradigm – program development through natural language interaction with users – to HPC programming, but also enables autonomous performance optimization with minimal user intervention through a sophisticated multi-agent design. To achieve these objectives, VibeCodeHPC implements three core functionalities: (1) configuration capabilities tailored to the unique development environments of supercomputers, (2) collaborative operation among multiple LLM agents with distinct roles – Project Manager (PM), System Engineer (SE), Programmer (PG), and Continuous Deliverer (CD), and (3) long-term autonomous operation through agent activity monitoring and dynamic deployment mechanisms. This paper highlights one of the most powerful features of VibeCodeHPC: fully automated code optimization through autonomous operation without user intervention. Specifically, it demonstrates the performance optimization of CPU-based codes on GPU-equipped systems for matrix multiplication and a Poisson equation solver using Jacobi’s iterative method. The results show that the multi-agent configuration employed in VibeCodeHPC enables faster and more reliable development of higher-performance code compared to a single-agent setup.
I Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities in code generation, automatically transforming natural language specifications into executable programs. These models, trained on large-scale source code corpora and natural language texts, have been applied to various aspects of software development, including rapid prototyping, code refactoring, and automated testing. More recently, coding agent systems have emerged that autonomously accomplish tasks initially specified by users through automatic iterative operations. Representative systems include Claude Code [^1] https://github.com/anthropics/claude-code, Codex CLI [^2] https://github.com/openai/codex, and Gemini CLI [^3] https://github.com/google-gemini/gemini-cli. These systems have become even more user-friendly tools for developers through integration with command-line interfaces (CLI). However, while LLM-based systems have shown promising results in generating functionally correct code, their application to performance-critical domains such as high-performance computing (HPC) applications executed on supercomputers remains limited.
HPC programming involves stringent requirements beyond functional correctness. For example, it requires the application of architecture-specific optimization techniques, achieving high performance particularly through thread and process parallelization, minimizing power consumption, and balancing numerical computation precision with computational speed. Additionally, HPC software typically utilizes parallel programming models such as MPI (Message Passing Interface), OpenMP, CUDA, and OpenACC. These models must be appropriately selected based on application characteristics and the target execution environment. Furthermore, utilizing supercomputer systems requires system-specific operations that differ significantly from program development on typical PCs. Specifically, this includes operating batch job systems, managing software modules, and managing Secure Shell (SSH) connections.
The complex development tasks described above pose difficult challenges not only for humans but also for code generation AI agents. Similar to humans, LLMs have memory capacity constraints (context windows), and when a single agent performs complex tasks for extended periods, situations may arise where it forgets initially given instructions. In such circumstances, the effectiveness of multi-agent systems where multiple LLMs with specialized roles operate collaboratively is expected. By having each agent focus on a limited role, it may be possible to reduce the memory burden on individual agents and reliably accomplish more complex tasks. This effect is considered particularly pronounced in highly complex development such as HPC programming. Moreover, in complex development such as HPC programming, creating rigorous requirement definitions in advance is also a difficult task for humans. Recently, a development paradigm called Vibe Coding [19][28] has been proposed, where programs are incrementally developed through dialogue between humans and LLMs without relying on strict requirement definitions. The term “Vibe” in Vibe Coding carries connotations of “getting into the groove” and “feeling the atmosphere,” with dialogue between humans and LLMs occurring as if it were informal communication. Through this dialogue, programs are gradually shaped. Such a development paradigm may be particularly effective for complex development targets where creating rigorous requirement definitions is difficult, such as HPC programming.
Based on the above background, this study proposes VibeCodeHPC, a multi-agent system based on LLMs for the automatic tuning of HPC programs on supercomputers. VibeCodeHPC adopts Claude Code as its backend and provides an integrated environment that facilitates program development in supercomputer settings. Thanks to its sophisticated multi-agent mechanism, the system not only enables vibe coding as if working with a “specialized HPC programming team,” but also achieves fully automated tuning through autonomous operation that minimizes user intervention. To realize these objectives, VibeCodeHPC provides the following core functionalities:
  1. 1.
    Configuration capabilities tailored to the unique development environments of supercomputers
  2. 2.
    Collaborative operation among multiple LLM agents with distinct roles – Project Manager (PM), System Engineer (SE), Programmer (PG), and Continuous Deliverer (CD)
  3. 3.
    Long-term autonomous operation through agent activity monitoring and dynamic deployment mechanisms
The structure of this paper is as follows. Section II reviews related work. Section III presents an overview of VibeCodeHPC and its implementation, with a particular focus on its sophisticated multi-agent features. Section IV demonstrates one of the most powerful features and challenges of VibeCodeHPC: fully automated code optimization through autonomous operation without user intervention. Specifically, it evaluates the performance optimization of CPU-based codes on GPU-equipped systems for matrix multiplication and a Poisson equation solver using Jacobi’s iterative method. Finally, Section V summarizes the findings of this study and discusses future research directions.
II Related Work
II-A LLM-Based Code Generation
Large language models (LLMs) including GPT-4, LLaMA 3, and Code LLaMA have significantly advanced automatic code generation. These models have demonstrated capabilities in producing syntactically correct and semantically meaningful code from natural language descriptions, with applications in software prototyping, automated bug fixing, and code refactoring. Chen et al. [4] showed that models trained on large-scale code corpora can generate effective solutions for programming contest problems. Austin et al. [2] and Li et al. [26] introduced prompt engineering and in-context learning strategies to improve code correctness and robustness. However, these studies primarily focus on general-purpose programming tasks without addressing HPC-specific performance constraints.
II-B HPC Code Generation and Optimization
HPC code generation traditionally relies on domain experts to manually implement performance-critical kernels using parallel programming models such as MPI, OpenMP, CUDA, and OpenACC. Compiler-assisted approaches [43] [23] have aimed to automate parts of this process by applying loop transformations, memory layout optimizations, and vectorization. More recent studies have investigated domain-specific languages (DSLs) [40] [16] to generate optimized HPC kernels from high-level specifications. While these methods can yield high performance, they typically require explicit user annotations or DSL expertise, limiting their accessibility. The integration of LLMs into HPC code generation remains largely unexplored, especially in conjunction with automated performance tuning.
LLM-based HPC code generation. Recent studies have begun evaluating LLMs for parallel programming models and HPC kernels, including OpenMP-centric models and MPI+OpenMP code generation in scalable workloads [3] [7]. Early evaluations of LLMs for OpenMP/MPI/CUDA/OpenACC kernels [10] [38] and LLM-assisted portability from Fortran to C/C++ [37] suggest that HPC-specific constraints substantially affect correctness and performance outcomes. These works provide concrete baselines against which VibeCodeHPC can be positioned.
HPC operations and applied studies. Practical case studies that combine LLM-based coding agents with GPU acceleration and testing pipelines offer direct evidence of feasibility in HPC contexts [14]. Additionally, HPC-oriented LLM systems and evaluations for software development provide a broader landscape for integration and operational support [46, 8, 11, 30].
II-C AT and Iterative Refinement
AT has been widely adopted in HPC to search parameter spaces and identify optimal configurations for target architectures. Frameworks such as ATLAS [42], OpenTuner [1], and GPTune [27] automate this process, often leveraging search algorithms like Bayesian optimization, genetic algorithms, or multi-armed bandits. Additionally, iterative refinement strategies—where code is repeatedly modified based on performance feedback—have been explored in machine learning-guided compilation [5] [12] . While these approaches are effective in optimizing existing code, their integration with LLM-driven code synthesis is still in its infancy.
Agentic optimization vs autotuning. Comparative analyses that pit agentic LLM approaches against classical autotuning for loop transformations suggest that LLM-driven strategies can complement or accelerate parameter search in certain settings [34]. This line of work motivates combining structured tuning with multi-agent orchestration.
II-D Iterative Prompt in LLM-based Code Generation
Recent studies have explored iterative prompting in code generation AI. DiCuffa et al. [6] used the DevGPT dataset to demonstrate that structural prompt patterns enable high-quality outputs with fewer iterations. Sorokin et al. [36] introduced an iterative self-learning loop with re-ranking models and approximate policy optimization, achieving performance surpassing GPT-4. Ye et al. [45] proposed Prochemy, which automatically refines initial prompts based on execution results, enhancing performance in code translation and HumanEval tasks.
However, iteration can introduce vulnerabilities. Shukla et al. [35] reported a 37.6% increase in critical vulnerabilities after five iterations, highlighting the need for human oversight. From a structural perspective, Ren et al. [32] proposed Knowledge-driven Prompt Chaining (KPC) for exception handling, while Li et al. [25] introduced Structured CoT, improving Pass@1 accuracy by up to 13.79% on HumanEval. Ridnik et al. [33] demonstrated that test-driven, multi-stage iterative flows improved Pass@5 accuracy from 19% to 44%.
Wei et al. [41] investigated CUDA code optimization using multi-agent systems without additional training. However, the role-sharing approach proposed in our study has not been evaluated for cross-language optimization.
These studies constitute methodological investigations into iterative prompting. Research on AI-agent collaboration paradigms, as exemplified by VibeCodeHPC, remains nascent.
Tech trends. Beyond research papers, the recent surge of general-purpose multi-agent frameworks (e.g., AutoGen[44], MetaGPT[13], CAMEL[24], and Magentic-One[9]) and agent orchestration tools (Claude Code subagents [^4] https://docs.anthropic.com/en/docs/claude-code/sub-agents, SuperClaude [^5] https://github.com/NomenAK/SuperClaude, Claude-Flow [^6] https://github.com/ruvnet/claude-flow, Claude Squad [^7] https://github.com/smtg-ai/claude-squad, Claude Task Master [^8] https://github.com/eyaltoledano/claude-task-master, Vibe Kanban [^9] https://github.com/BloopAI/vibe-kanban, KAMUI [^10] https://3rd.kamui.ai/en) highlights the growing interest in collaborative LLM workflows. Microsoft’s agent orchestration patterns categorize Magentic orchestration as a distinct mode, which aligns closely with VibeCodeHPC’s collaborative, role-separated workflow; however, a framework explicitly specialized for supercomputer code development and operations remains unproposed. [29].
II-E Multi-agent for HPC Code Generation
Recent HPC-focused multi-agent systems have begun to target code generation, testing, and optimization workflows. Examples include multi-agent HPC code generation and optimization [31], MPI code synthesis with LLMs [39], and multi-agent unit test generation for HPC applications [18]. Profiling-guided autonomous code generation for parallel workloads is also emerging as a closely related direction [17]. These approaches highlight the growing feasibility of collaborative agentic workflows in HPC and motivate VibeCodeHPC’s emphasis on role separation, monitoring, and dynamic agent deployment.
II-F Summary
Existing research in LLM-based code generation has achieved significant progress in producing functionally correct programs, but it rarely addresses HPC-specific performance optimization. On the other hand, the HPC community has developed powerful auto-tuning and compiler-based optimization techniques, yet these methods often lack the flexibility and adaptability of LLM-based code synthesis. Bridging these domains—by integrating LLM-driven generation, HPC performance tuning, and iterative improvement—remains an open challenge, which our proposed VibeCodeHPC framework seeks to address.
III Overview and Implementation of VibeCodeHPC
This section explains the system overview of VibeCodeHPC, the operational flow from the user’s perspective, and the implementation that realizes it. In particular, we provide detailed explanation of the multi-agent implementation, which is the core innovation of VibeCodeHPC.
III-A System Components
VibeCodeHPC consists of the following three main components.
  • CLI-based Coding AI Agent: This component is the LLM agent that actually performs autonomous programming. The current VibeCodeHPC adopts Claude Code, but conceptually it is also possible to adopt other CLI coding agents. Claude Code is an LLM-based coding assistance tool developed by Anthropic that operates in CLI environments. It receives instructions from users and has the ability to autonomously perform code generation, editing, execution, debugging, and other tasks.
  • Functionalities for Supercomputer Operation and Development: CLI coding agents such as Claude Code are designed with general-purpose programming applications in mind. Therefore, additional support is required when handling operations and functions specific to supercomputers. Accordingly, VibeCodeHPC provides the following supercomputer-specific functionalities:
    • SSH connection management: Establishment and maintenance of remote access to supercomputers (Figure 1)
    • Batch job system operation: Job submission, execution status monitoring, and result retrieval
    • Software module management: Loading compilers and libraries using the environment module command
    • Hardware information provision: System information such as cluster configuration and GPU availability
    • Software environment information provision: Information about available compilers, libraries, etc.
    This information is provided to Claude Code to support agents in appropriately performing system-specific operations.
  • Multi-agent Collaborative Operation Functionality: This is the core functionality provided by VibeCodeHPC. Since coding agents such as Claude Code operate as single-agent systems by default, various implementations are necessary to enable collaborative operation of multiple agents. This component includes the following functionalities:
    • Agent configuration and role setting: Definition of each agent’s responsibility scope and operational rules
    • Inter-agent communication mechanism: Message transmission and reception between agents
    • Multi-agent operation rules: Control rules for collaborative operation
    • Agent operation monitoring functionality: Real-time monitoring of each agent’s activity status
    • Dynamic agent deployment functionality: Addition and removal of agents as needed
    Details of this component are described in Section III-C.
Figure 1: SSH connection architecture. Private keys are managed by ssh-agent and not exposed to the LLM context. Desktop Commander MCP server maintains persistent SSH/SFTP sessions.
III-B Workflow
The typical usage flow of VibeCodeHPC is shown below.
  1. 1.
    System Startup: VibeCodeHPC can be operated anywhere, but considering the typical usage pattern of supercomputers, it is assumed that VibeCodeHPC operates on the user’s local terminal (Linux environment) that accesses the supercomputer. The user starts VibeCodeHPC on the local terminal, and VibeCodeHPC launches Claude Code.
  2. 2.
    Provision of Tasks and Requirement Definitions: Users provide VibeCodeHPC with the code they want to optimize for performance and, if necessary, a requirement definition document. When utilizing the Vibe Coding development paradigm, a requirement definition document is not mandatory, and users can proceed with work incrementally while interacting with the system. However, when autonomous code optimization without human intervention is desired, it is recommended to prepare the requirement definition document as rigorously as possible. VibeCodeHPC provides a template for requirement definition documents targeting general supercomputers to support user creation. Items to be described in the requirement definition document include:
    • Target hardware: CPU architecture, GPU availability, node configuration, etc.
    • Performance goals: Target execution time, throughput, energy efficiency, etc.
    • Task constraints: Parts prohibited from optimization, prohibited libraries, precision requirements, etc.
    The more specific these descriptions are, the higher the likelihood of improved LLM work accuracy.
  3. 3.
    Execution of Autonomous Optimization: When a requirement definition document is provided, VibeCodeHPC begins the optimization process. The code generation [^11] The code generation LLM itself runs on Anthropic’s cloud. is performed on the local side, and the generated code is then transferred to the supercomputer via the Secure File Transfer Protocol (SFTP). Subsequently, compilation and execution are carried out on the supercomputer through SSH. For efficient management of SSH and SFTP sessions, VibeCodeHPC employs Desktop Commander MCP [^12] https://github.com/wonderwhy-er/DesktopCommanderMCP, which is particularly recommended for supercomputers that require two-factor authentication. Job execution is managed by batch job schedulers such as PBS Pro or Slurm. The entire workflow is orchestrated by Claude Code, while VibeCodeHPC provides the corresponding operational instructions. VibeCodeHPC operates continuously until the user-specified goal is achieved, but users can intervene at any time if necessary. When adopting the Vibe Coding development paradigm, users can appropriately interact with VibeCodeHPC and incrementally refine the optimization direction.
  4. 4.
    Monitoring and Result Confirmation: Users can monitor the operation of each agent in real time. Additionally, VibeCodeHPC automatically generates work logs, allowing users to grasp optimization progress by reviewing them. When VibeCodeHPC achieves the user’s objectives, a final report is automatically generated. This report includes the optimization process, achieved performance, techniques used, etc.
III-C Multi-agent Mechanism
For multiple agents to operate collaboratively and perform productive activities as a multi-agent system, various technical challenges must be overcome. Particularly in programming for supercomputers, problems arise that cannot be addressed by general multi-agent systems due to the complexity of their operation and programming. VibeCodeHPC implements various innovations to overcome these challenges.
III-C1 Agent Configuration
VibeCodeHPC enables collaborative operation of multiple agents specialized for different tasks. Specifically, the following agents are defined. The roles of these agents are defined by markdown-format configuration files that Claude Code loads at startup, and users can customize them.
  • Project Manager (PM): The top-level agent that supervises the entire project. Based on user instructions, it manages the project, issues instructions to subordinate agents, monitors their operations, and also serves as the dialogue interface with users.
  • System Engineer (SE): An intermediate management layer agent that manages the work of Programmers (PG) based on PM’s instructions. It coordinates the work of multiple programmers and collects statistical information and generates reports.
  • Programmer (PG): An agent responsible for actual programming tasks. PGs are deployed for each specialized field. Examples of specialized fields include CPU optimization, GPU optimization (CUDA, OpenACC), distributed parallelization using MPI, and algorithm optimization. This deployment has templates prepared according to the system configuration of the target supercomputer. For example, for GPU-equipped systems, CUDA-specialized PGs and OpenACC-specialized PGs are deployed.
  • Continuous Deliverer (CD): An agent responsible for version control of generated code, pushing to GitHub, document generation, etc. It manages and publishes development deliverables continuously.
Figure 2 shows the directory structure used in the multi-agent configuration, illustrating how PM, SE, and PGs organize their workspaces.
Figure 2: Project directory structure in multi-agent configuration. PM spawns SE and CD, with PGs specialized for different parallelization techniques (OpenMP, MPI, CUDA).
III-C2 Inter-agent Communication Mechanism
VibeCodeHPC utilizes a terminal multiplexer (terminal software that can manage multiple sessions) called Tmux to enable user monitoring of agent activities and communication between agents. Tmux provides functionality to simultaneously display multiple sessions divided into panes on a GUI and message sending functionality between sessions (tmux send-keys command). VibeCodeHPC utilizes “Tmux Multi-agent Communication,” [^13] https://github.com/nishimoto265/Claude-Code-Communication an inter-agent communication library for Claude Code that leverages this functionality. What communication each agent performs with other agents is described in the markdown configuration file that defines each agent’s operation, for example, ”PG must report execution results to PM”, ”When errors occur, report to SE and request guidance” and so on. These communication rules enable information sharing and collaborative operation between agents. Figure 3 shows an example of the desktop during multi-agent operation, where each agent runs in a separate tmux pane.
Figure 3: Desktop view during multi-agent operation with tmux panes
III-C3 Agent Management and Dynamic Deployment by PM
When the system starts, the PM agent launches first and loads the markdown configuration file describing the PM’s role. The PM launches other agents according to its contents and manages subsequent operations. In existing multi-agent implementations (e.g., Tmux Multi-agent Communication Demo), a method was adopted that fixedly assigns agent configurations at startup. However, VibeCodeHPC implements a more flexible dynamic agent deployment functionality. The dynamic agent deployment functionality allows the PM to dynamically change the number and specialization of PGs according to situations such as:
  • Addition of specialized agents according to requirement definition contents
  • Adjustment according to target system configuration (CPU/GPU, parallelization model, etc.)
  • Resource reallocation according to work progress status
This functionality allows the system to flexibly respond to diverse tasks.
III-C4 Guaranteeing Persistent Agent Operation
LLM-based agents operate according to initially given instructions, but when the LLM’s context window (memory area) becomes full, problems arise where initial instructions are overlooked. In the case of Claude Code, when the context window becomes full, an automatic compression function called ”auto-compact” operates. This function compresses old information to secure memory capacity, but compression is not perfect. As a result, cases may occur where agents forget initial instructions or cause unexpected behavior leading to activity cessation. To address such situations, VibeCodeHPC implements the following mechanisms:
  • Recovery Processing Using Hooks Functionality: Recovery processing is implemented for when agents stop operating using the hooks functionality provided by Claude Code. Hooks are a function that automatically executes pre-prepared commands when specific events (e.g., agent stoppage) occur. In VibeCodeHPC, when an agent stops, hooks execute processing to reload the initial instruction contents. This allows agents to return to their initial state and guarantees persistent operation.
  • Context Usage Status Monitoring: To prevent the occurrence of such unexpected situations, VibeCodeHPC monitors each agent’s context usage status in real time. Based on these monitoring results, when the context of other agents approaches fullness, the PM agent issues instructions to those agents to reconfirm initial instructions. This enables prevention of information loss due to auto-compact. Regarding PM’s own context management, we leverage the characteristic that PM, being the top-level agent and not engaging in frequent dialogue, consumes context relatively slowly.
  • Visualization of Context Usage Status: Each agent’s context usage status is visualized in graphs, allowing users to grasp it in real time. This visualization functionality enables users to intervene as needed and adjust system operation. Figure 4 shows an example of context visualization for each agent in the matrix multiplication optimization experiment described later.
Figure 4: Example of context visualization for multi-agent configuration
III-D Advantages of VibeCodeHPC
Through the above design and implementation, VibeCodeHPC provides the following advantages:
  1. 1.
    Long-term Autonomous Operation: The persistent agent operation guarantee mechanism enables continuation of optimization work for extended periods without human intervention.
  2. 2.
    Flexible Task Response: Dynamic agent deployment enables flexible response to diverse system configurations and task requirements.
  3. 3.
    Enhanced Robustness: Mutual monitoring between agents and recovery mechanisms realize higher robustness compared to single-agent systems.
  4. 4.
    Support for Vibe Coding Paradigm: Supporting both autonomous operation and user intervention enables accommodation of both rigorous requirement definitions and informal dialogue.
IV Demonstration
This section demonstrates the major contribution of VibeCodeHPC: code optimization through autonomous multi-agent behavior. As a case study, we optimize the CPU-targeted naive matrix multiplication code and the Himeno benchmark [^14] https://i.riken.jp/en/supercom/documents/himenobmt/ – which is the kernel code of a Poisson equation solver using Jacobi’s iterative method for computational fluid dynamics (CFD) computation – for a GPU-equipped machine.
IV-A Experimental Setup
We describe the experimental setup commonly used for matrix-multiplication and Himeno benchmark cases.
We conducted experiments using the Supercomputer “Flow” Type II Subsystem [15] at the Information Technology Center, Nagoya University. The system specifications include: two Intel Xeon Gold 6230 CPUs (20 cores, 2.10–3.90 GHz); four NVIDIA Tesla V100 GPUs (2560 FP64 cores, up to 1530 MHz); 384 GiB DDR4 main memory (2933 MHz); and 32 GiB HBM2 device memory per GPU. The theoretical peak performance reaches 33.888 TFLOPS in double precision (1.344 TFLOPS per CPU socket, 7.8 TFLOPS per GPU). Memory bandwidth comprises 281.5 GB/s for main memory and 900 GB/s per GPU for device memory. GPU interconnection uses NVLINK2 (50 GB/s bidirectional), while CPU-GPU connection employs PCI-Express 3.0 (x16). GCC 4.8.5 and CUDA 11.2 are used for the compilation of generated codes. The following software is in use on the local side for VibeCodeHPC.
  • VibeCodeHPC-jp v0.7.5151515https://github.com/Katagiri-Hoshino-Lab/VibeCodeHPC-jp/releases/tag/v0.7.5
  • Windows 11 (local PC)
    • WSL version: 2.4.13.0
    • Kernel Version: 5.15.167.4-1
    • Ubuntu 22.04.5 LTS (Jammy Jellyfish)
    • tmux 3.2a
    • git version 2.34.1
    • gh version 2.4.0+dfsg1 (2022-03-23 Ubuntu 2.4.0+dfsg1-2)
  • OpenSSH_8.9p1 Ubuntu-3ubuntu0.13, OpenSSL 3.0.2 15 Mar 2022
  • nvm 0.39.7
  • Node.js v22.16.0
  • npm 10.9.2
  • Claude Code 2.0.8 – Claude Sonnet 4.5
  • Desktop Commander MCP v0.2.17
  • jq-1.6
  • Python 3.10.12
    • pip 22.0.2 from /usr/lib/python3/dist-packages/pip (python 3.10)
    • matplotlib 3.10.5
    • numpy 2.2.6
    • pandas 2.3.2
    • scipy 1.15.3
The agent configuration consists of one PM, one SE, four PGs, and one CD. The role of each are described in Section III-C. PGs are specialized and deployed for each parallelization technique. For example, in EX2 of matrix multiplication, PG1 handled CUDA, PG2 handled OpenMP, PG3 handled MPI, and PG4 handled OpenACC optimization. The evaluation compared using these multi-agents (“Multi”) with having a single agent perform the same roles (“Solo”). Due to the random nature of Claude Code’s prompt responses and generated code, three experiments were conducted under identical conditions for each case (EX1, EX2, EX3).
The following common items were specified in the requirement definition document for both experiments. The optimization goal was set to maximize throughput, with the instruction to “continue optimization to the limit and aim for the highest performance on a single node.” The budget for computing resource consumption (utilization points based on job execution time and resource usage) was specified as: minimum consumption: 100 points, target: 500 points, upper limit: 1000 points. Figure 5 shows an example of budget tracking visualization from the matrix multiplication multi-agent EX3 experiment, automatically generated by parsing the ChangeLog.md maintained by agents. The following supplementary instructions were given to each agent: CD should frequently push all code except system information to GitHub; SE should exclude code that does not meet requirements when creating graphs; all agents should record both performance (GFLOPS) and execution time, with GFLOPS as the unified vertical axis. Note that the reported performance values include CPU-GPU data transfer time.
Figure 5: Budget tracking visualization from matrix multiplication EX3. Shows actual consumption (blue), prediction (purple dashed), and threshold levels: minimum 100pt (green), target 500pt (orange), upper limit 1000pt (red).
IV-B Matrix multiplication
Matrix multiplication is one of the fundamental kernels in scientific computing. This computation is highly compute-intensive; therefore, even when accounting for the data transfer cost between the CPU and GPU, utilizing the GPU – with its high theoretical computational performance – is expected to be advantageous.
We provided a naive C-language implementation using a triple-nested loop and optimized it using VibeCodeHPC. The code performs double-precision matrix multiplication on row-major matrices, with the i, j, and k loops arranged from the outermost to innermost levels. An OpenMP parallel for directive is applied to the outermost i loop. A separate program is provided for verifying correctness and measuring performance. This program is used solely for performance evaluation and is excluded from the optimization target. The problem size is fixed at m = n = k = 2048. Functionality is verified by comparing the results with those of OpenBLAS, and correctness is confirmed when the maximum relative error for random input is less than or equal to 1e-8. Performance is reported in GFLOPS and milliseconds.
The following constraints were specified in the requirement definition document for the matrix multiplication experiment:
  1. 1.
    Time limit: minimum 45 minutes, maximum 60 minutes
  2. 2.
    Job class: cx-single (one node of Type II)
  3. 3.
    Parallelization techniques: OpenMP, MPI, CUDA, OpenACC, etc. may be freely used. However, the use of numerical computation libraries such as cuBLAS or MKL is prohibited
Table I shows the performance of the generated code. The technology classification in the tables (OpenMP, MPI, SIMD, CUDA, OpenACC, etc.) corresponds to the directory names where PGs were deployed, representing the parallelization technique each PG handled. Empty cells in the tables indicate that the corresponding parallelization technique was not selected by the agent. In each experiment, the choice of which parallelization technique to use was left to the autonomous judgment of the agents. We can see that the multi-agent configuration achieves the maximum performance in almost all cases. Table II shows the number of generated codes. We see that the multi-agent configuration generates more variations.
Notably, the solo agent once ignored the user’s instructions, submitted code using cuBLAS, and completed optimization without noticing. In the multi-agent configuration, such rule violations were monitored by other agents and automatically detected, preventing them.
Additionally, in EX3, the solo agent violated the requirement definition that fixed the matrix size at 2048, gradually expanding the matrix size to 16384 to improve performance, and ultimately reported achieving 8898 GFLOPS.
TABLE I: Performance of optimized matrix-multiplication code (in GFLOPS)
Cat Technology Mode Experiments
EX1 EX2 EX3 Avg
CPU OpenMP Multi 197.7 213.0 189.6 200.1
Solo
MPI Multi 6.7 30.7 18.7
Solo
SIMD Multi 183.5 183.5
Solo
GPU CUDA Multi 1384.6 3677.2 916.2 1993.0
Solo 804.5 606.9 1097.7 837.0
OpenACC Multi 249.8 249.8
Solo
Hybrid MPI+CUDA Multi 332.1 332.1
Solo
TABLE II: The number of generated codes (matrix multiplication)
Cat Technology Mode Experiments
EX1 EX2 EX3 Total
CPU OpenMP Multi 13 4 8 25
Solo
MPI Multi 3 4 0 7
Solo
SIMD Multi 0 8 0 8
Solo
GPU CUDA Multi 7 7 13 27
Solo 7 5 6 18
OpenACC Multi 11 0 0 11
Solo
Hybrid MPI+CUDA Multi 0 0 6 6
Solo
IV-C Himeno Benchmark (Poisson equation solver)
The Himeno Benchmark is a CFD benchmark developed by RIKEN to evaluate memory bandwidth performance. It is widely used for comparing the effective performance of supercomputers. It simulates an iterative solver for the three-dimensional Poisson equation. Specifically, it performs a simple iterative computation similar to the Jacobi method, using a 7-point difference approximation via the finite difference method (FDM).
We targeted the “himenoBTM.c” in “Himeno Benchmark 98.” It includes several problem size settings, and we selected SMALL among them. This is too small to offload to the GPU, so CPU utilization is expected. The code includes a problem generation section containing embedded fixed values. To prevent these from being expanded into the computation section and thus defeating the code’s purpose, random perturbations are added to the matrix generation. These perturbations are configured to have no impact on the operation verification. The code displays performance (in GFLOPS) and verifies the correctness of the computation at the end. The correctness is judged when the “gosa” item is 3e-3 or less and the maximum error of all coefficient values is 1e-6 or less.
The following constraints were specified in the requirement definition document for the Himeno benchmark experiment:
  1. 1.
    Time limit: minimum 60 minutes, maximum 90 minutes
  2. 2.
    Job class: cx-share (1/4 of one Type II node, i.e., one V100 and 10-core CPU)
  3. 3.
    Parallelization techniques: Try various patterns including OpenMP, MPI, CUDA, OpenACC, etc.
Table III shows the achieved performance. This case is unsuitable for GPU implementation, and CPU-oriented optimization achieved good performance. Table IV shows the number of generated codes. Figure 6 shows the best performance progression chart created by the SE agent during the multi-agent Himeno benchmark optimization experiment EX3.
Figure 6: Example of best performance progression chart created by SE during multi-agent Himeno benchmark optimization experiment
TABLE III: Performance of optimized Himeno benchmark code (in GFLOPS)
Cat Tech Mode Experiments
EX1 EX2 EX3 Avg
CPU OpenMP Multi 47.2 77.8 67.1 64.03
Solo 21.5 23.9 39.9 28.43
MPI Multi 40.1 57.6 61.1 52.93
Solo
OpenMP+MPI Multi 68.8 54.2 61.5
Solo
GPU OpenACC Multi 24.4 20.6 20.3 21.77
Solo 20.5 20.5
CUDA Multi 21.1 24.5 23.3 22.97
Solo 14.1 19.5 16.8
TABLE IV: The number of generated codes (Himeno benchmark)
Cat Technology Mode Experiments
EX1 EX2 EX3 Total
CPU OpenMP Multi 8 3 14 25
Solo 2 4 3 9
MPI Multi 5 8 6 19
Solo
OpenMP+MPI Multi 6 1 0 7
Solo
GPU OpenACC Multi 10 5 2 17
Solo 1 0 0 1
CUDA Multi 8 13 9 30
Solo 3 0 2 5
In this experiment, a problem was discovered where the CUDA code generated for the Himeno benchmark lacked synchronization for parallel processing. This problem was not detected during the experiments because it passed the correctness verification tests provided by the user. Such problems are difficult to detect with test cases because they depend on the timing of parallel processing. It is important to provide LLMs with knowledge that appropriate synchronization processing and double-buffer implementation are necessary for parallelization such as iterative solvers.
IV-D Discussion
The experimental results confirmed the effectiveness of the multi-agent configuration from multiple perspectives. A recent study suggests that multi-agent performance is sensitive to coordination topology and task structure, and gains are not uniform across configurations, motivating topology-aware orchestration. [47]
Task Execution Stability: Pushing to GitHub and the creation of best performance progression graphs were more stable with the multi-agent configuration. In the solo agent experiments, pushing was not performed in 5 out of 6 experiments, and the best performance progression graphs had insufficient data points.
Prompt Compliance: The solo agent tended to be satisfied with its own optimization before the specified “minimum time” (e.g., at the 30-minute mark from the start), created a final report, and stopped generating code thereafter. Once it declared the project finished, even when the requirement definition was re-presented through hooks with instructions such as “try various patterns” and “continue optimization to the limit,” it repeatedly responded “the project is finished” and often repeated sleep commands without effectively using the given time limit. One contributing factor is that the solo agent failed to track whether it had reached the given budget. Although we provided Python code that can track supercomputer point consumption in real time by parsing job start/end times recorded in ChangeLog.md in a specified format, there were cases where the agent forgot about its existence or failed to track consumption due to missing entries in the ChangeLog, resulting in 0 points being displayed.
Performance Superiority: In EX1 and EX2 of matrix multiplication, despite the similar number of CUDA code generations between multi and solo (Table II), multi achieved significantly higher performance. The reason why solo reached lower final performance is considered to be that prompts and execution logs related to Git operations and visualization, which are not directly related to code generation, consumed context and hindered concentration on code optimization. In the multi-agent configuration, these ancillary tasks are distributed to specialized agents such as CD and SE, enabling PG to focus on code optimization.
Performance Degradation in EX3: On the other hand, in EX3 of matrix multiplication, multi’s performance was lower than in EX1 and EX2. This is considered to be partly due to PM adopting a different directory structure than in EX1 and EX2. While EX1 and EX2 followed a policy of gradually transitioning from single technologies (OpenMP, CUDA, etc.) to hybrid technologies (MPI+CUDA, etc.), EX3 created directories for hybrid technologies from the beginning. Although three experiments are insufficient for statistical certainty, this suggests the effectiveness of a gradual approach.
Quality Assurance through Mutual Monitoring: As mentioned in the matrix multiplication experiment, in cases where the solo agent ignored user instructions and used cuBLAS, the multi-agent configuration automatically detected and prevented this through monitoring by other agents. This mutual monitoring function plays an important role in long-term autonomous operation.
V Conclusion
This study proposed “VibeCodeHPC,” a multi-LLM agent system aimed at automatically tuning HPC programs on supercomputers. This system is designed to specialize a CLI coding agent (Claude Code in this study) for complex HPC programming development by configuring multiple agents with distinct roles that autonomously coordinate their actions. Consequently, VibeCodeHPC not only enables vibe coding with a specialized HPC programming team but also achieves complete automatic tuning through autonomous operation that minimizes user intervention. This paper demonstrates VibeCodeHPC’s most challenging feature – fully autonomous performance optimization – using CPU-oriented matrix multiplication for GPU systems and a Poisson equation solver using Jacobi’s iterative method. This evaluation shows that the multi-agent configuration effectively monitors each other, enabling the generation of more accurate code that closely follows user requirements in a shorter time.
VibeCodeHPC offers a wide range of features, and we hope each one will be evaluated for its effectiveness. However, we wish to emphasize that we have demonstrated its most challenging feature—autonomous operation—which inherently encompasses the evaluation of the effectiveness of its other features. For example, while VibeCodeHPC aims for vibe coding, it allows user intervention at any time through the tmux screen, enabling the incorporation of vibe coding development techniques whenever needed. The advanced autonomous operation achieved by VibeCodeHPC can accelerate development through vibe coding, and with user intervention, VibeCodeHPC can more easily develop the desired program. Furthermore, developing distributed parallel programs is crucial on supercomputers. However, VibeCodeHPC already possesses the capability to handle batch systems and MPI, addressing this within the same fundamental framework as single-node GPU programming. MPI programming capability depends on the underlying Claude Code’s capabilities, but VibeCodeHPC provides the supporting environment for this.
VibeCodeHPC has many directions for expansion. For example, since VibeCodeHPC is an automatic tuning system, integrating existing automatic tuning strategies or tools [20][21][22] with it is expected to enable more efficient automatic tuning. Furthermore, in developing numerical computation code, unlike programs where mere functionality suffices, it is necessary to address challenges specific to numerical computation, particularly numerical error issues. Additionally, mixed-precision approaches are effective for acceleration, necessitating the expansion of features that incorporate such optimization considerations.
The introduction of automatic verification methods such as dependency verification through data flow analysis is also being considered, but there are practical constraints such as execution time limitations. In VibeCodeHPC, strengthening the code-level verification function by audit agents to automatically discover problems that are difficult to detect with user-provided tests alone is a future challenge.
The VibeCodeHPC code is available [^16] https://github.com/Katagiri-Hoshino-Lab/VibeCodeHPC-jp.
Acknowledgment
This work was supported by the Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures (JHPCN) and the High-Performance Computing Infrastructure (HPCI) under Project ID jh250015. It was also partially supported by JSPS KAKENHI Grant Numbers JP23K11126 and JP24K02945. In addition, this work was supported by the JST Research and Development Program for Next-generation Edge AI Semiconductors (Grant Number JPMJES2511).

### v3

@@
 I Introduction
-In recent years, large language models (LLMs) have demonstrated remarkable capabilities in code generation, automatically transforming natural language specifications into executable programs.
-These models, trained on large-scale source code corpora and natural language texts, have been applied to various aspects of software development, including rapid prototyping, code refactoring, and automated testing.
-More recently, coding agent systems have emerged that autonomously accomplish tasks initially specified by users through automatic iterative operations.
-Representative systems include Claude Code [^1] https://github.com/anthropics/claude-code, Codex CLI [^2] https://github.com/openai/codex, and Gemini CLI [^3] https://github.com/google-gemini/gemini-cli.
-These systems have become even more user-friendly tools for developers through integration with command-line interfaces (CLI).
-However, while LLM-based systems have shown promising results in generating functionally correct code, their application to performance-critical domains such as high-performance computing (HPC) applications executed on supercomputers remains limited.
-HPC programming involves stringent requirements beyond functional correctness.
-For example, it requires the application of architecture-specific optimization techniques, achieving high performance particularly through thread and process parallelization, minimizing power consumption, and balancing numerical computation precision with computational speed.
-Additionally, HPC software typically utilizes parallel programming models such as MPI (Message Passing Interface), OpenMP, CUDA, and OpenACC.
-These models must be appropriately selected based on application characteristics and the target execution environment.
-Furthermore, utilizing supercomputer systems requires system-specific operations that differ significantly from program development on typical PCs.
-Specifically, this includes operating batch job systems, managing software modules, and managing Secure Shell (SSH) connections.
-The complex development tasks described above pose difficult challenges not only for humans but also for code generation AI agents.
-Similar to humans, LLMs have memory capacity constraints (context windows), and when a single agent performs complex tasks for extended periods, situations may arise where it forgets initially given instructions.
-In such circumstances, the effectiveness of multi-agent systems where multiple LLMs with specialized roles operate collaboratively is expected.
-By having each agent focus on a limited role, it may be possible to reduce the memory burden on individual agents and reliably accomplish more complex tasks.
-This effect is considered particularly pronounced in highly complex development such as HPC programming.
-Moreover, in complex development such as HPC programming, creating rigorous requirement definitions in advance is also a difficult task for humans.
-Recently, a development paradigm called Vibe Coding [19][28] has been proposed, where programs are incrementally developed through dialogue between humans and LLMs without relying on strict requirement definitions.
-The term “Vibe” in Vibe Coding carries connotations of “getting into the groove” and “feeling the atmosphere,” with dialogue between humans and LLMs occurring as if it were informal communication.
-Through this dialogue, programs are gradually shaped.
-Such a development paradigm may be particularly effective for complex development targets where creating rigorous requirement definitions is difficult, such as HPC programming.
+In recent years, large language models (LLMs) have shown strong capabilities in automatic code generation, transforming natural language specifications into executable programs.
+Trained on large-scale source code corpora, these models can generate syntactically correct and semantically meaningful code and have been applied to software development tasks such as prototyping, bug fixing, refactoring, and testing.
+Prior work has demonstrated that large-scale code-trained models can generate effective solutions for programming problems [4], and that prompt engineering and in-context learning improve code correctness and robustness [2][26].
+More recently, coding agent systems have emerged that autonomously perform user-specified tasks through iterative execution and refinement.
+Representative examples include Claude Code [^1] https://github.com/anthropics/claude-code, Codex CLI [^2] https://github.com/openai/codex, and Gemini CLI [^3] https://github.com/google-gemini/gemini-cli, which integrate LLMs with command-line interfaces (CLI).
+However, despite these advances, existing LLM-based systems primarily target general-purpose programming and do not sufficiently address performance-critical domains such as high-performance computing (HPC) applications on supercomputers, particularly with respect to performance optimization.
+HPC programming imposes stringent requirements beyond functional correctness, including architecture-specific optimizations, parallelization across threads and processes, and trade-offs between numerical precision, performance, and power consumption.
+HPC applications typically rely on parallel programming models such as MPI, OpenMP, CUDA, and OpenACC, which must be selected according to application characteristics and execution environments.
+In addition, supercomputer usage requires system-specific operations, including batch job execution, software module management, and Secure Shell (SSH) access.
+These complexities pose significant challenges for both human developers and code generation AI agents.
+LLMs, like humans, are constrained by limited context windows, and prolonged complex tasks may lead to forgotten instructions or degraded behavior.
+Multi-agent systems, in which multiple LLMs with specialized roles collaborate, offer a promising approach to mitigating these limitations by distributing cognitive load across agents.
+This advantage is particularly relevant for complex domains such as HPC programming.
+Moreover, defining rigorous requirements in advance is often difficult in such settings.
+To address this, the Vibe Coding paradigm has been proposed [19][28], in which programs are incrementally developed through interactive dialogue between humans and LLMs without strict upfront specifications.
+This paradigm is well suited to complex development domains, including HPC programming.
 Based on the above background, this study proposes VibeCodeHPC, a multi-agent system based on LLMs for the automatic tuning of HPC programs on supercomputers.
-VibeCodeHPC adopts Claude Code as its backend and provides an integrated environment that facilitates program development in supercomputer settings.
+VibeCodeHPC is built on top of Claude Code and offers an integrated framework for program development targeting supercomputing environments.
 Thanks to its sophisticated multi-agent mechanism, the system not only enables vibe coding as if working with a “specialized HPC programming team,” but also achieves fully automated tuning through autonomous operation that minimizes user intervention.
@@
     Long-term autonomous operation through agent activity monitoring and dynamic deployment mechanisms
-The structure of this paper is as follows.
+This paper is organized as follows.
 Section II reviews related work.
-Section III presents an overview of VibeCodeHPC and its implementation, with a particular focus on its sophisticated multi-agent features.
-Section IV demonstrates one of the most powerful features and challenges of VibeCodeHPC: fully automated code optimization through autonomous operation without user intervention.
-Specifically, it evaluates the performance optimization of CPU-based codes on GPU-equipped systems for matrix multiplication and a Poisson equation solver using Jacobi’s iterative method.
-Finally, Section V summarizes the findings of this study and discusses future research directions.
+Section III presents VibeCodeHPC and its implementation, focusing on multi-agent features.
+Section IV demonstrates automated code optimization through autonomous operation.
+Finally, Section V summarizes the findings and discusses future research directions.
 II Related Work
-II-A LLM-Based Code Generation
-Large language models (LLMs) including GPT-4, LLaMA 3, and Code LLaMA have significantly advanced automatic code generation.
-These models have demonstrated capabilities in producing syntactically correct and semantically meaningful code from natural language descriptions, with applications in software prototyping, automated bug fixing, and code refactoring.
-Chen et al.
-[4] showed that models trained on large-scale code corpora can generate effective solutions for programming contest problems.
-Austin et al.
-[2] and Li et al.
-[26] introduced prompt engineering and in-context learning strategies to improve code correctness and robustness.
-However, these studies primarily focus on general-purpose programming tasks without addressing HPC-specific performance constraints.
-II-B HPC Code Generation and Optimization
-HPC code generation traditionally relies on domain experts to manually implement performance-critical kernels using parallel programming models such as MPI, OpenMP, CUDA, and OpenACC.
-Compiler-assisted approaches [43] [23] have aimed to automate parts of this process by applying loop transformations, memory layout optimizations, and vectorization.
-More recent studies have investigated domain-specific languages (DSLs) [40] [16] to generate optimized HPC kernels from high-level specifications.
-While these methods can yield high performance, they typically require explicit user annotations or DSL expertise, limiting their accessibility.
-The integration of LLMs into HPC code generation remains largely unexplored, especially in conjunction with automated performance tuning.
-LLM-based HPC code generation.
+II-A HPC Code Generation and Optimization
+HPC code generation has traditionally relied on domain experts to manually implement performance-critical kernels using parallel programming models such as MPI, OpenMP, CUDA, and OpenACC.
+Compiler-assisted approaches [43][23] partially automate this process through loop transformations, memory layout optimization, and vectorization, while domain-specific languages (DSLs) [40][16] generate optimized kernels from high-level specifications.
+Although these methods can achieve high performance, they often require explicit user annotations or DSL expertise, limiting accessibility.
 Recent studies have begun evaluating LLMs for parallel programming models and HPC kernels, including OpenMP-centric models and MPI+OpenMP code generation in scalable workloads [3] [7].
@@
 These works provide concrete baselines against which VibeCodeHPC can be positioned.
-HPC operations and applied studies.
 Practical case studies that combine LLM-based coding agents with GPU acceleration and testing pipelines offer direct evidence of feasibility in HPC contexts [14].
 Additionally, HPC-oriented LLM systems and evaluations for software development provide a broader landscape for integration and operational support [46, 8, 11, 30].
-II-C AT and Iterative Refinement
-AT has been widely adopted in HPC to search parameter spaces and identify optimal configurations for target architectures.
-Frameworks such as ATLAS [42], OpenTuner [1], and GPTune [27] automate this process, often leveraging search algorithms like Bayesian optimization, genetic algorithms, or multi-armed bandits.
-Additionally, iterative refinement strategies—where code is repeatedly modified based on performance feedback—have been explored in machine learning-guided compilation [5] [12] .
-While these approaches are effective in optimizing existing code, their integration with LLM-driven code synthesis is still in its infancy.
-Agentic optimization vs autotuning.
-Comparative analyses that pit agentic LLM approaches against classical autotuning for loop transformations suggest that LLM-driven strategies can complement or accelerate parameter search in certain settings [34].
-This line of work motivates combining structured tuning with multi-agent orchestration.
-II-D Iterative Prompt in LLM-based Code Generation
+II-B Auto-tuning and Iterative Refinement
+Auto-tuning (AT) has been widely adopted in HPC to explore parameter spaces and identify optimal configurations for target architectures.
+Frameworks such as ATLAS [42], OpenTuner [1], and GPTune [27] automate this process using search algorithms including Bayesian optimization, genetic algorithms, and multi-armed bandits.
+In addition, iterative refinement strategies, in which code is repeatedly modified based on performance feedback, have been studied in machine learning–guided compilation [5] [12].
+While these approaches effectively optimize existing code, their integration with LLM-driven code synthesis remains limited.
+Comparative studies contrasting agentic LLM approaches with classical autotuning for loop transformations indicate that LLM-driven methods can complement or accelerate parameter search in certain settings [34].
+These findings motivate the combination of structured tuning with multi-agent orchestration.
+II-C Iterative Prompt in LLM-based Code Generation
 Recent studies have explored iterative prompting in code generation AI.
@@
 However, the role-sharing approach proposed in our study has not been evaluated for cross-language optimization.
-These studies constitute methodological investigations into iterative prompting. Research on AI-agent collaboration paradigms, as exemplified by VibeCodeHPC, remains nascent.
-Tech trends.
-Beyond research papers, the recent surge of general-purpose multi-agent frameworks (e.g., AutoGen[44], MetaGPT[13], CAMEL[24], and Magentic-One[9]) and agent orchestration tools (Claude Code subagents [^4] https://docs.anthropic.com/en/docs/claude-code/sub-agents, SuperClaude [^5] https://github.com/NomenAK/SuperClaude, Claude-Flow [^6] https://github.com/ruvnet/claude-flow, Claude Squad [^7] https://github.com/smtg-ai/claude-squad, Claude Task Master [^8] https://github.com/eyaltoledano/claude-task-master, Vibe Kanban [^9] https://github.com/BloopAI/vibe-kanban, KAMUI [^10] https://3rd.kamui.ai/en) highlights the growing interest in collaborative LLM workflows.
-Microsoft’s agent orchestration patterns categorize Magentic orchestration as a distinct mode, which aligns closely with VibeCodeHPC’s collaborative, role-separated workflow; however, a framework explicitly specialized for supercomputer code development and operations remains unproposed.
-[29].
+These studies constitute methodological investigations into iterative prompting.
+Research on AI-agent collaboration paradigms, as exemplified by VibeCodeHPC, remains nascent.
+II-D Multi-Agent Frameworks and Orchestration Tools
+Beyond research papers, the recent surge of general-purpose multi-agent frameworks (e.g., AutoGen[44], MetaGPT[13], CAMEL[24], and Magentic-One[9]) and agent orchestration tools (Claude Code subagents [^4] https://docs.anthropic.com/en/docs/claude-code/sub-agents, SuperClaude [^5] https://github.com/NomenAK/SuperClaude, Claude-Flow [^6] https://github.com/ruvnet/claude-flow, Claude Squad [^7] https://github.com/smtg-ai/claude-squad, Vibe Kanban [^8] https://github.com/BloopAI/vibe-kanban, KAMUI [^9] https://3rd.kamui.ai/en) highlights the growing interest in collaborative LLM workflows.
+Microsoft’s agent orchestration patterns categorize Magentic orchestration as a distinct mode, which aligns closely with VibeCodeHPC’s collaborative, role-separated workflow; however, a framework specialized for supercomputer code development remains unproposed[29].
 II-E Multi-agent for HPC Code Generation
@@
 II-F Summary
-Existing research in LLM-based code generation has achieved significant progress in producing functionally correct programs, but it rarely addresses HPC-specific performance optimization.
-On the other hand, the HPC community has developed powerful auto-tuning and compiler-based optimization techniques, yet these methods often lack the flexibility and adaptability of LLM-based code synthesis.
-Bridging these domains—by integrating LLM-driven generation, HPC performance tuning, and iterative improvement—remains an open challenge, which our proposed VibeCodeHPC framework seeks to address.
+Existing research in LLM-based code generation has achieved progress in producing functionally correct programs, but it rarely addresses HPC-specific performance optimization.
+On the other hand, the HPC community has developed powerful auto-tuning and compiler-based optimization techniques, yet these methods often lack the flexibility of LLM-based code synthesis.
+Bridging these domains – by integrating LLM-driven generation, HPC performance tuning, and iterative improvement—remains an open challenge, which our proposed VibeCodeHPC framework seeks to address.
+Figure 1: Workflow of VibeCodeHPC.
 III Overview and Implementation of VibeCodeHPC
-This section explains the system overview of VibeCodeHPC, the operational flow from the user’s perspective, and the implementation that realizes it.
-In particular, we provide detailed explanation of the multi-agent implementation, which is the core innovation of VibeCodeHPC.
+This section explains the system overview of VibeCodeHPC, the operational flow from the user’s perspective, and the implementation realizing it.
+In particular, we provide explanation of the multi-agent implementation, which is the core innovation of VibeCodeHPC.
 III-A System Components
-VibeCodeHPC consists of the following three main components.
+VibeCodeHPC consists of the following main components.
   • CLI-based Coding AI Agent: This component is the LLM agent that actually performs autonomous programming.
@@
 Accordingly, VibeCodeHPC provides the following supercomputer-specific functionalities:
-    • SSH connection management: Establishment and maintenance of remote access to supercomputers (Figure 1)
+    • SSH connection management: Establishment and maintenance of remote access to supercomputers
     • Batch job system operation: Job submission, execution status monitoring, and result retrieval
@@
     • Software environment information provision: Information about available compilers, libraries, etc.
-    This information is provided to Claude Code to support agents in appropriately performing system-specific operations.
+    This information is provided to support agents in appropriately performing system-specific operations.
   • Multi-agent Collaborative Operation Functionality: This is the core functionality provided by VibeCodeHPC.
@@
     Details of this component are described in Section III-C.
-Figure 1: SSH connection architecture. Private keys are managed by ssh-agent and not exposed to the LLM context. Desktop Commander MCP server maintains persistent SSH/SFTP sessions.
 III-B Workflow
-The typical usage flow of VibeCodeHPC is shown below.
+Figure 1 illustrates the workflow of VibeCodeHPC, which is described as follows.
   1. 1.
-    System Startup: VibeCodeHPC can be operated anywhere, but considering the typical usage pattern of supercomputers, it is assumed that VibeCodeHPC operates on the user’s local terminal (Linux environment) that accesses the supercomputer.
-The user starts VibeCodeHPC on the local terminal, and VibeCodeHPC launches Claude Code.
+    System Startup: Although VibeCodeHPC can be operated anywhere, considering the usage pattern of supercomputers, it is assumed to run on the user’s local Linux terminal that accesses the supercomputer, from which it launches Claude Code.
   2. 2.
@@
   4. 4.
-    Monitoring and Result Confirmation: Users can monitor the operation of each agent in real time.
-Additionally, VibeCodeHPC automatically generates work logs, allowing users to grasp optimization progress by reviewing them.
-When VibeCodeHPC achieves the user’s objectives, a final report is automatically generated.
-This report includes the optimization process, achieved performance, techniques used, etc.
+    Monitoring and Result Confirmation: Users can monitor each agent’s operation in real time, and VibeCodeHPC automatically generates work logs that allow users to track optimization progress.
+When the user’s objectives are achieved, a final report is generated, summarizing the optimization process, achieved performance, and applied techniques.
 III-C Multi-agent Mechanism
@@
 VibeCodeHPC enables collaborative operation of multiple agents specialized for different tasks.
-Specifically, the following agents are defined.
-The roles of these agents are defined by markdown-format configuration files that Claude Code loads at startup, and users can customize them.
+Their roles are defined in markdown-format configuration files loaded by Claude Code at startup and can be customized.
   • Project Manager (PM): The top-level agent that supervises the entire project.
@@
   • Programmer (PG): An agent responsible for actual programming tasks.
-PGs are deployed for each specialized field.
-Examples of specialized fields include CPU optimization, GPU optimization (CUDA, OpenACC), distributed parallelization using MPI, and algorithm optimization.
-This deployment has templates prepared according to the system configuration of the target supercomputer.
-For example, for GPU-equipped systems, CUDA-specialized PGs and OpenACC-specialized PGs are deployed.
+PGs are deployed for each specialized field, such as CPU optimization, GPU optimization (CUDA, OpenACC), distributed parallelization using MPI, and algorithm optimization.
+Deployment templates are prepared according to the system configuration of the target supercomputer; for example, GPU-equipped systems deploy CUDA- and OpenACC-specialized PGs.
   • Continuous Deliverer (CD): An agent responsible for version control of generated code, pushing to GitHub, document generation, etc. It manages and publishes development deliverables continuously.
-Figure 2 shows the directory structure used in the multi-agent configuration, illustrating how PM, SE, and PGs organize their workspaces.
+Figure 2 shows the directory structure of the multi-agent configuration, illustrating the organization of PM, SE, and PG workspaces.
 Figure 2: Project directory structure in multi-agent configuration. PM spawns SE and CD, with PGs specialized for different parallelization techniques (OpenMP, MPI, CUDA).
@@
 VibeCodeHPC utilizes “Tmux Multi-agent Communication,” [^12] https://github.com/nishimoto265/Claude-Code-Communication an inter-agent communication library for Claude Code that leverages this functionality.
-What communication each agent performs with other agents is described in the markdown configuration file that defines each agent’s operation, for example, ”PG must report execution results to PM”, ”When errors occur, report to SE and request guidance” and so on.
+What communication each agent performs with other agents is described in the markdown configuration file that defines each agent’s operation, for example, “PG must report execution results to PM,” “When errors occur, report to SE and request guidance” and so on.
 These communication rules enable information sharing and collaborative operation between agents.
 Figure 3 shows an example of the desktop during multi-agent operation, where each agent runs in a separate tmux pane.
-Figure 3: Desktop view during multi-agent operation with tmux panes
+Figure 3: Desktop view of multi-agent operation with tmux panes
 III-C3 Agent Management and Dynamic Deployment by PM
-When the system starts, the PM agent launches first and loads the markdown configuration file describing the PM’s role.
-The PM launches other agents according to its contents and manages subsequent operations.
-In existing multi-agent implementations (e.g., Tmux Multi-agent Communication Demo), a method was adopted that fixedly assigns agent configurations at startup.
-However, VibeCodeHPC implements a more flexible dynamic agent deployment functionality.
-The dynamic agent deployment functionality allows the PM to dynamically change the number and specialization of PGs according to situations such as:
+At system startup, the PM agent is launched first and loads a markdown configuration file describing its role.
+Based on this configuration, the PM launches other agents and manages subsequent operations.
+Existing multi-agent implementations (e.g., Tmux Multi-agent Communication Demo) typically fix agent configurations at startup, whereas VibeCodeHPC supports flexible dynamic agent deployment.
+This functionality allows the PM to dynamically adjust the number and specialization of PGs according to situations such as:
   • Addition of specialized agents according to requirement definition contents
@@
 III-C4 Guaranteeing Persistent Agent Operation
-LLM-based agents operate according to initially given instructions, but when the LLM’s context window (memory area) becomes full, problems arise where initial instructions are overlooked.
-In the case of Claude Code, when the context window becomes full, an automatic compression function called ”auto-compact” operates.
-This function compresses old information to secure memory capacity, but compression is not perfect.
-As a result, cases may occur where agents forget initial instructions or cause unexpected behavior leading to activity cessation.
-To address such situations, VibeCodeHPC implements the following mechanisms:
+LLM-based agents operate according to initially given instructions; however, when the LLM’s context window becomes full, these instructions may be overlooked.
+In Claude Code, an automatic compression mechanism called auto-compact compresses older information to free memory, but this process is imperfect and may lead to forgotten instructions or unexpected behavior, including activity cessation.
+To address these issues, VibeCodeHPC implements the following mechanisms:
   • Recovery Processing Using Hooks Functionality: Recovery processing is implemented for when agents stop operating using the hooks functionality provided by Claude Code.
@@
 GCC 4.8.5 and CUDA 11.2 are used for the compilation of generated codes.
-The following software is in use on the local side for VibeCodeHPC.
-  • VibeCodeHPC-jp v0.7.5151515https://github.com/Katagiri-Hoshino-Lab/VibeCodeHPC-jp/releases/tag/v0.7.5
-  • Windows 11 (local PC)
-    • WSL version: 2.4.13.0
-    • Kernel Version: 5.15.167.4-1
-    • Ubuntu 22.04.5 LTS (Jammy Jellyfish)
-    • tmux 3.2a
-    • git version 2.34.1
-    • gh version 2.4.0+dfsg1 (2022-03-23 Ubuntu 2.4.0+dfsg1-2)
-  • OpenSSH_8.9p1 Ubuntu-3ubuntu0.13, OpenSSL 3.0.2 15 Mar 2022
-  • nvm 0.39.7
-  • Node.js v22.16.0
-  • npm 10.9.2
-  • Claude Code 2.0.8 – Claude Sonnet 4.5
-  • Desktop Commander MCP v0.2.17
-  • jq-1.6
-  • Python 3.10.12
-    • pip 22.0.2 from /usr/lib/python3/dist-packages/pip (python 3.10)
-    • matplotlib 3.10.5
-    • numpy 2.2.6
-    • pandas 2.3.2
-    • scipy 1.15.3
+The local environment for VibeCodeHPC consists of: VibeCodeHPC-jp v0.7.5141414https://github.com/Katagiri-Hoshino-Lab/VibeCodeHPC-jp/releases/tag/v0.7.5, Windows 11 with WSL 2.4.13.0 (Ubuntu 22.04.5 LTS, tmux 3.2a, git 2.34.1, gh 2.4.0), OpenSSH 8.9p1, Node.js v22.16.0, npm 10.9.2, Claude Code 2.0.8 (Claude Sonnet 4.5), Desktop Commander MCP v0.2.17, jq 1.6, and Python 3.10.12 (matplotlib 3.10.5, numpy 2.2.6, pandas 2.3.2, scipy 1.15.3).
 The agent configuration consists of one PM, one SE, four PGs, and one CD.
@@
 The following common items were specified in the requirement definition document for both experiments.
-The optimization goal was set to maximize throughput, with the instruction to “continue optimization to the limit and aim for the highest performance on a single node.” The budget for computing resource consumption (utilization points based on job execution time and resource usage) was specified as: minimum consumption: 100 points, target: 500 points, upper limit: 1000 points.
-Figure 5 shows an example of budget tracking visualization from the matrix multiplication multi-agent EX3 experiment, automatically generated by parsing the ChangeLog.md maintained by agents.
-The following supplementary instructions were given to each agent: CD should frequently push all code except system information to GitHub; SE should exclude code that does not meet requirements when creating graphs; all agents should record both performance (GFLOPS) and execution time, with GFLOPS as the unified vertical axis.
-Note that the reported performance values include CPU-GPU data transfer time.
+The optimization goal was to maximize throughput, with instructions to continue optimization to the limit and achieve the highest possible single-node performance.
+The computing resource budget (utilization points based on job execution time and resource usage) was defined as 100 points minimum, 500 points target, and 1000 points upper limit.
+Figure 5 shows an example of budget tracking visualization from the matrix multiplication experiment (Multi EX3), automatically generated by parsing the ChangeLog.md maintained by agents.
+Supplementary instructions were also provided: CD should frequently push all non-system code to GitHub; SE should exclude non-compliant code when generating graphs; and all agents should record both performance and execution time, using GFLOPS as the unified vertical axis.
+Reported performance values include CPU–GPU data transfer time.
 Figure 5: Budget tracking visualization from matrix multiplication EX3.
 Shows actual consumption (blue), prediction (purple dashed), and threshold levels: minimum 100pt (green), target 500pt (orange), upper limit 1000pt (red).
+Figure 6: Best performance progression for matrix multiplication across all experiments.
+Multi-agent runs (solid lines) generally achieve higher GFLOPS than solo-agent runs (dashed lines) within the 60-minute time limit.
 IV-B Matrix multiplication
-Matrix multiplication is one of the fundamental kernels in scientific computing.
-This computation is highly compute-intensive; therefore, even when accounting for the data transfer cost between the CPU and GPU, utilizing the GPU – with its high theoretical computational performance – is expected to be advantageous.
+Matrix multiplication is a fundamental kernel in scientific computing.
+This computation is highly compute-intensive; therefore, even accounting for data transfer costs between the CPU and GPU, utilizing GPUs with high theoretical computational performance is expected to be advantageous.
 We provided a naive C-language implementation using a triple-nested loop and optimized it using VibeCodeHPC.
@@
 We see that the multi-agent configuration generates more variations.
+Figure 6 shows the best performance progression over time for all experiments in both multi-agent and solo configurations.
 Notably, the solo agent once ignored the user’s instructions, submitted code using cuBLAS, and completed optimization without noticing.
@@
 Additionally, in EX3, the solo agent violated the requirement definition that fixed the matrix size at 2048, gradually expanding the matrix size to 16384 to improve performance, and ultimately reported achieving 8898 GFLOPS.
+A similar matrix size violation also occurred in the multi-agent EX2, where the CUDA PG expanded the matrix size to 4096 and 8192 to improve arithmetic intensity, achieving up to 3677 GFLOPS.
+Although these results were reported in the inter-agent communication logs, neither the SE nor PM agents detected the deviation from the fixed problem size.
+The performance values in Table I exclude results obtained with non-standard matrix sizes for both cases.
 TABLE I: Performance of optimized matrix-multiplication code (in GFLOPS)
-Cat Technology Mode Experiments
+Target Technology Mode Experiments
 EX1 EX2 EX3 Avg
-CPU OpenMP Multi 197.7 213.0 189.6 200.1
+CPU OpenMP Multi 105.3 213.0 189.6 169.3
 Solo
@@
 Solo
-GPU CUDA Multi 1384.6 3677.2 916.2 1993.0
-Solo 804.5 606.9 1097.7 837.0
+GPU CUDA Multi 1384.6 1141.0 916.2 1147.3
+Solo 804.5 561.7 1097.7 821.3
 OpenACC Multi 249.8 249.8
@@
 TABLE II: The number of generated codes (matrix multiplication)
-Cat Technology Mode Experiments
+Target Technology Mode Experiments
 EX1 EX2 EX3 Total
@@
 Specifically, it performs a simple iterative computation similar to the Jacobi method, using a 7-point difference approximation via the finite difference method (FDM).
-We targeted the “himenoBTM.c” in “Himeno Benchmark 98.” It includes several problem size settings, and we selected SMALL among them.
-This is too small to offload to the GPU, so CPU utilization is expected.
-The code includes a problem generation section containing embedded fixed values.
-To prevent these from being expanded into the computation section and thus defeating the code’s purpose, random perturbations are added to the matrix generation.
-These perturbations are configured to have no impact on the operation verification.
-The code displays performance (in GFLOPS) and verifies the correctness of the computation at the end.
-The correctness is judged when the “gosa” item is 3e-3 or less and the maximum error of all coefficient values is 1e-6 or less.
+We targeted the “himenoBTM.c” in “Himeno Benchmark 98.” It includes several problem size settings, and we selected SMALL ($129\times 65\times 65$ grid, approximately 29 MB of single-precision working data across seven arrays—three multi-component coefficient matrices, pressure, boundary, and two work arrays) with its default iteration count ($\mathrm{NN}=200$).
+The sub-second execution time enables both configurations to complete many compile–test–evaluate cycles within the time budget.
+At this scale, both CPU-targeted and GPU-based optimizations are potentially effective, requiring agents to evaluate diverse parallelization strategies—complementing the matrix multiplication experiment where GPU is clearly dominant.
+This is a conservative setting: the fast turnaround gives the solo agent ample optimization opportunities, whereas larger problem sizes would reduce cycle counts and widen the multi-agent advantage.
+Evaluation at larger scales remains future work.
+The original code contains embedded fixed coefficient values (zeros and ones).
+Without modification, a compiler or an LLM-generated code could apply constant propagation, eliminating the memory-access patterns that the benchmark is designed to measure and producing artificially inflated performance.
+To prevent this, we add small random perturbations to the coefficient matrices; these perturbations are bounded so as not to affect correctness verification.
+The code reports performance in GFLOPS and verifies correctness: the residual norm “gosa” must satisfy $\mathrm{gosa}\leq 3\times 10^{-3}$, and the maximum coefficient error must be at most $10^{-6}$.
 The following constraints were specified in the requirement definition document for the Himeno benchmark experiment:
@@
 Table III shows the achieved performance.
-This case is unsuitable for GPU implementation, and CPU-oriented optimization achieved good performance.
+As expected from the small problem size, CPU-targeted optimizations (OpenMP and MPI) significantly outperformed GPU-based implementations.
+Notably, multi-agent configurations explored diverse parallelization strategies and achieved the best results via hybrid OpenMP+MPI, whereas solo agents produced fewer non-GPU codes.
 Table IV shows the number of generated codes.
-Figure 6 shows the best performance progression chart created by the SE agent during the multi-agent Himeno benchmark optimization experiment EX3.
-Figure 6: Example of best performance progression chart created by SE during multi-agent Himeno benchmark optimization experiment
+Figure 7 shows the best performance progression over time for all Himeno benchmark experiments in both multi-agent and solo configurations.
+Figure 7: Best performance progression for the Himeno benchmark across all experiments.
+Multi-agent runs (solid lines) consistently outperform solo-agent runs (dashed lines) within the 90-minute time limit.
 TABLE III: Performance of optimized Himeno benchmark code (in GFLOPS)
-Cat Tech Mode Experiments
+Target Tech Mode Experiments
 EX1 EX2 EX3 Avg
@@
 Solo 20.5 20.5
-CUDA Multi 21.1 24.5 23.3 22.97
+CUDA Multi 21.1 24.5 8.9 18.17
 Solo 14.1 19.5 16.8
 TABLE IV: The number of generated codes (Himeno benchmark)
-Cat Technology Mode Experiments
+Target Technology Mode Experiments
 EX1 EX2 EX3 Total
@@
 [47]
-Task Execution Stability: Pushing to GitHub and the creation of best performance progression graphs were more stable with the multi-agent configuration.
+Task Execution Stability: Pushing to GitHub and the creation of best performance progression graphs were more stable with the multi-agent configuration, as shown in Figures 6 and 7.
 In the solo agent experiments, pushing was not performed in 5 out of 6 experiments, and the best performance progression graphs had insufficient data points.
@@
 This mutual monitoring function plays an important role in long-term autonomous operation.
+However, this monitoring was not comprehensive: in EX2, the CUDA PG’s expansion of the matrix size beyond the specified $n=2048$ went undetected, despite being visible in the ChangeLog and inter-agent communication logs.
+This suggests that the current monitoring mechanism is more effective at detecting explicit rule violations, such as prohibited library usage, than implicit constraint deviations, such as problem size changes.
 V Conclusion
-This study proposed “VibeCodeHPC,” a multi-LLM agent system aimed at automatically tuning HPC programs on supercomputers.
-This system is designed to specialize a CLI coding agent (Claude Code in this study) for complex HPC programming development by configuring multiple agents with distinct roles that autonomously coordinate their actions.
-Consequently, VibeCodeHPC not only enables vibe coding with a specialized HPC programming team but also achieves complete automatic tuning through autonomous operation that minimizes user intervention.
-This paper demonstrates VibeCodeHPC’s most challenging feature – fully autonomous performance optimization – using CPU-oriented matrix multiplication for GPU systems and a Poisson equation solver using Jacobi’s iterative method.
-This evaluation shows that the multi-agent configuration effectively monitors each other, enabling the generation of more accurate code that closely follows user requirements in a shorter time.
-VibeCodeHPC offers a wide range of features, and we hope each one will be evaluated for its effectiveness.
-However, we wish to emphasize that we have demonstrated its most challenging feature—autonomous operation—which inherently encompasses the evaluation of the effectiveness of its other features.
-For example, while VibeCodeHPC aims for vibe coding, it allows user intervention at any time through the tmux screen, enabling the incorporation of vibe coding development techniques whenever needed.
-The advanced autonomous operation achieved by VibeCodeHPC can accelerate development through vibe coding, and with user intervention, VibeCodeHPC can more easily develop the desired program.
-Furthermore, developing distributed parallel programs is crucial on supercomputers.
-However, VibeCodeHPC already possesses the capability to handle batch systems and MPI, addressing this within the same fundamental framework as single-node GPU programming.
-MPI programming capability depends on the underlying Claude Code’s capabilities, but VibeCodeHPC provides the supporting environment for this.
-VibeCodeHPC has many directions for expansion.
-For example, since VibeCodeHPC is an automatic tuning system, integrating existing automatic tuning strategies or tools [20][21][22] with it is expected to enable more efficient automatic tuning.
-Furthermore, in developing numerical computation code, unlike programs where mere functionality suffices, it is necessary to address challenges specific to numerical computation, particularly numerical error issues.
-Additionally, mixed-precision approaches are effective for acceleration, necessitating the expansion of features that incorporate such optimization considerations.
-The introduction of automatic verification methods such as dependency verification through data flow analysis is also being considered, but there are practical constraints such as execution time limitations.
-In VibeCodeHPC, strengthening the code-level verification function by audit agents to automatically discover problems that are difficult to detect with user-provided tests alone is a future challenge.
+This study proposed VibeCodeHPC, a multi-LLM agent system for automatic tuning of HPC programs on supercomputers.
+By orchestrating multiple role-specialized agents around a CLI coding agent, VibeCodeHPC enables vibe coding for complex HPC development and achieves fully automated tuning with minimal user intervention.
+This paper demonstrates this capability using a CPU-targeted matrix multiplication implementation and a Poisson equation solver based on Jacobi’s iterative method for GPU systems, and the evaluation showed that mutual monitoring among agents improves adherence to user requirements and reduces development time.
+While VibeCodeHPC provides a wide range of features, this study focuses on demonstrating autonomous operation, the most challenging aspect of the system.
+Because autonomous operation requires effective integration of other functionalities, its evaluation reflects their combined effectiveness.
+VibeCodeHPC also allows optional user intervention via the tmux screen, supporting interactive refinement when needed.
+Furthermore, VibeCodeHPC supports batch job systems and MPI within the same framework as single-node GPU programming, providing the execution and orchestration environment required for distributed parallel development.
+VibeCodeHPC has several directions for future expansion.
+As an automatic tuning system, integrating existing auto-tuning strategies and tools [20][21][22] is expected to further improve tuning efficiency.
+In addition, extending VibeCodeHPC to address numerical computation–specific challenges, such as numerical error control and mixed-precision optimization, is an important direction.
+Automatic verification methods, including dependency verification via data flow analysis, are also being considered, although practical constraints such as execution time remain.
+Strengthening code-level verification through audit agents to detect issues beyond the coverage of user-provided tests is therefore a key future challenge.
 The VibeCodeHPC code is available [^15] https://github.com/Katagiri-Hoshino-Lab/VibeCodeHPC-jp.